An element of R is called idempotent if a 2=a. Show that a division ring contains exactly 2 idempotent elements.



.



1
Expert's answer
2021-09-06T18:00:35-0400

Suppose "a\in R" is such that "a^2=a". We have two possibilities :

  • "a=0", as "0\cdot 0 = 0",
  • "a\neq 0", in which case there exists "a^{-1}" and thus we have "id=a\cdot a^{-1}=(a\cdot a)\cdot a^{-1} = a\cdot id = a", so "a=id".

In addition, as in a ring we have "id \neq 0" by definition, then the set of idempotent elements consists of 2 elements : and "id".